OpenStudy (anonymous):

Solve: −2(2x + 5) − 3 = −3(x − 1)

OpenStudy (anonymous):

First, distribute -2 across (2x+5). That means multiply -2 by both the 2x and by 5. Rewrite the left side after you do that. Do the same thing with the -3 on the right. OK so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-4-10-3=-3x-3

OpenStudy (anonymous):

Right, except that the first term would be -4x. Now combine like terms on the left and right sides.

OpenStudy (anonymous):

Did you get -4x-13=-3x+3? (It should be +3 on the right side because you multiplied -3*-1 when you distributed.)

OpenStudy (anonymous):

Terrell - are you there? We're almost done!

OpenStudy (anonymous):

I have to go now - here is a quick look at the remaining steps: -4x-13=-3x+3 Add 4x to both sides of the equation Subtract 3 from both sides of the equation That's it. You should then have the answer. I'll check back later in case you have any questions.